Circular No. 8933

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2008 B1–B4, C/2008 C3–C5, AND C/2008 E4 (SOHO)

Further to IAUC 8932, additional apparent comets have been found on SOHO website images; all are Kreutz sungrazers except for C/2008 C3 (no known group) and C/2008 E4 (Kracht group). In the tabulated column for instrument below, an asterisk (*) indicates that astrometry was also measured and reduced from HI1 images aboard the STEREO-A spacecraft. C/2008 B1 was tiny, stellar, and faint (mag ∼ 7). C/2008 B2 reached mag ∼ 6 in H1A and C3 images and was just barely visible with a thin, faint tail in C2 images. C/2008 B3, C/2008 B4, and C/2008 C4 were very faint (mag 7.5), the first and third being slightly diffuse and the second being quite diffuse. C/2008 C3 was stellar in appearance and brightened slowly and gradually, reaching mag ∼ 7 as it left the field-of-view. C/2008 C5 was also stellar in appearance, reaching mag ∼ 6.5.

C/2008 E4 is described by K. Battams as “typically [for a Kracht-group object] small, stellar, and faint (maybe peaking at mag 7 as it entered the C2 coronagraph), fading fast as it neared the sun”. R. Kracht suggested the identity of the Kracht group comets C/2002 Q8 and C/2008 E4 on the basis of the MPEC 2002-Q46 measurements of the former and his own measurements of the latter; Kracht gave the previous perihelion date T = 1997 Feb. 14.9 but was unable to locate the comet on C2 images from that time. Linked orbital elements (P = 5.52 yr) for 2002 and 2008 by B. G. Marsden appear on MPEC 2008-F32.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 B1 Jan. 26.221 20^{h}46.^{m}1 −21^{o}03′ C3  HS 2008-G59
C/2008 B2  27.138 21 00.0 −21 43 C3/2* AK 2008-G59
C/2008 B3  28.696 20 57.9 −19 43 C3  HS 2008-G59
C/2008 B4  29.904 21 03.8 −19 33 C3  BZ 2008-G59
C/2008 C3 Feb. 3.146 21 02.9 −18 12 C2  RK 2008-G60
C/2008 C4  3.821 21 22.3 −18 16 C3*  HS 2008-G60
C/2008 C5  5.738 21 33.9 −17 30 C3*  AK 2008-G60
C/2008 E4 Mar. 3.254 23 02.6 − 5 18 C2  HS 2008-F32

COMET C/2008 E1 (CATALINA) Prediscovery astrometry (MPEC 2008-F36) shows this comet (IAUC

8923) to be of intermediate period; elements from MPEC 2008-G33: epoch 2008 Aug. 2.0 TT, T = 2008 Aug. 11.284 TT, q = 4.82971 AU, e = 0.54759, ω = 269^{o}.910, Ω = 189^{o}.035, i = 35^{o}.038 (equinox 2000.0), P = 34.9 years.

2008 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green